I am stumped on how to write the following equation
Ernie has $3.65 in nickels and dimes. He has 13 more nickels than dimes. how many coins does he have?
Help

Found 2 solutions by funmath, checkley71:
Answer by funmath(2933)

Ernie has $.65 in nickels and dimes. He has 13 more nickels than dimes. how many coins does he have?
:
Let the number of dimes be: x
then the number of nickles is: x+13
Then the amount of money in dimes he has is: .10x
The amount of money in nickles is: .05(x+13)
The total amount of money is: $3.65
:
The problem to solve is:
.10x+.05(x+13)=3.65
.10x+.05x+.65=3.65
.15x+.65=3.65
.15x+.65-.65=3.65-.65
.15x=3.00
.15x/.15=3.00/.15
x=20
The amount of dimes: x=20
The amount of nickles: x+13=20+13=33
The total amount of coins are:20+33=53
:
Check does 20 dimes and 33 nickles = $3.65?
.10(20)+.05(33)=$3.65
2.00+1.65=$3.65
$3.65=$3.65 Cool beans, we're right!!!
Happy Calculating!!!

Answer by checkley71(8403)

5N+10D=65
N=D+13
5(D+13)+10D=65
5D+65+10D=65
15D=0 NO DIMES
5N+0=65
5N=65
N=65/5
N=13 NICKELS